OpenStudy (anonymous):

Which point is a solution to the following system of inequalities? x – 2y > 10 3x + 4y < 10

hero (hero):

I'm guessing you have a list of answer choices since there are many possible points that are included in the solution.

OpenStudy (anonymous):

(1, 10) (2, –3) (4, –4) (–2, –5)

OpenStudy (anonymous):

options, haha sorry

hero (hero):

Okay, so let me ask you this. You have the list of answer choices. What is preventing you from substituting each point into both equations to test and see if it works?

OpenStudy (anonymous):

I wanna know how to get to the answer properly

hero (hero):

Okay, so the list of points above are (x,y) points correct?

OpenStudy (anonymous):

yes

hero (hero):

Okay, so for example, the first possible point (1,10) means that x = 1 and y = 10, correct?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

go on..?

OpenStudy (radar):

@brittneyy18rm who is to say that testing the points are not the proper way to see if they satisfy the inequalities?

OpenStudy (radar):

What method have you been instructed to use? I would do just as Hero suggested.

OpenStudy (anonymous):

ahh alright that's true, I'll plug in the points then

OpenStudy (anonymous):

I got the 3rd one, but I think it's wrong

OpenStudy (radar):

There is a method to this madness. Take a look at the 2nd equation which in my mind is simpler than the first equation. 3x +4y < 10 lets examine the points listed and see if we can eliminate any. Then what we have left we will test on the first equation. A. (1,10) test: 3+40 <10 A. Is a no go. B. (2,-3) test 6 - 12 = -6 B. Is a possibilit C.(4, -4) Also a possibility D (-2, -5) Also a possibility Now to the first equation x -2y > 10 B. 2,-3) 2-2(-3) > 10, 8 > 10 No Go for B (we have already eliminated A) C. (4, -4) 4 - 2(-4)>10 12>10 A Go Looks like answer C but check out D for kicks D (-2,-5) -2 - 2(-5)>10 8>10 No Go

OpenStudy (radar):

The web was down for bit there sorry for delay.

OpenStudy (anonymous):

oh yay I was right haha

OpenStudy (radar):

Yes, good deal, good luck with these.

OpenStudy (anonymous):

would you mind helping me with my last question? it's kinda different: \[\sqrt[2]{x+7}^{-10}=-8\]

OpenStudy (anonymous):

options: x = –6 x = –4 x = 32 x = 74

OpenStudy (radar):

\[\sqrt{x+7^{-10}}=-8\]\[\sqrt{x+7}^{-10}=-8\] Which of these two is it?

OpenStudy (anonymous):

the 2nd one

OpenStudy (radar):

\[1 \over \sqrt{x+7})^{10}=-8\]\[1/(x+7)^{5}=-8\]

OpenStudy (radar):

I am having some trouble, are you suppose to be using logs?

OpenStudy (anonymous):

I'm not sure

OpenStudy (radar):

I am at a loss, I would have to guess at this one, and I would guess at x=-4, cause it doesn't make sense to me. I would repost that problem and someone with more knowledge than I can help you. Suggesting you guess is not a good idea at this time. Please post it as a new question.

OpenStudy (radar):

Are you sure that the radical has an index of 2? (square root)???